OpenStudy (anonymous):

Can someone help me with this fractions question? How do you get the left side to equal the right side?

OpenStudy (anonymous):

Ive attached a screen shot. Thanks!

OpenStudy (michele_laino):

hint: we can write: \[\large {x^6} + 3{x^4} + 3{x^2} + 1 = {\left( {{x^2} + 1} \right)^3}\]

OpenStudy (anonymous):

Yep... I can see that, but would the expression be \[-\frac{ x }{ (x^2+1)^3 }\] if you subtract?

OpenStudy (anonymous):

Sorry not cube just \[-\frac{ x }{ (x^2+1) }\]

OpenStudy (michele_laino):

what is the least common multiple between \[{\left( {{x^2} + 1} \right)^3}\] and \[{\left( {{x^2} + 1} \right)^2}\]?

OpenStudy (anonymous):

\[(x^2+1)^5\] ?

OpenStudy (michele_laino):

no, it is \[{\left( {{x^2} + 1} \right)^3}\]

OpenStudy (michele_laino):

so the right side can be rewrtten as below: \[\frac{x}{{{{\left( {{x^2} + 1} \right)}^2}}} - \frac{x}{{{{\left( {{x^2} + 1} \right)}^3}}} = \frac{{x\left( {{x^2} + 1} \right) - x \cdot 1}}{{{{\left( {{x^2} + 1} \right)}^3}}} = ...?\]

OpenStudy (anonymous):

OK! How did you make the numerator of the right side like that??? I get the answer now! Just thats confusing, how does x^3 become that whole thing :S

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

My brain shuts down after studying for exams...

OpenStudy (anonymous):

I get it now

OpenStudy (anonymous):

I hope they dont ask this on the test but if they do ill remember this haha thanks

OpenStudy (michele_laino):

the left side is: \[\frac{{{x^3}}}{{{{\left( {{x^2} + 1} \right)}^3}}}\] since: \[{x^6} + 3{x^4} + 3{x^2} + 1 = {\left( {{x^2} + 1} \right)^3}\]

OpenStudy (michele_laino):

thanks :)

OpenStudy (anonymous):

Yep I got it :) This is just the first step to the problem, I actually have to integrate that lol

OpenStudy (anonymous):

Sigh

OpenStudy (michele_laino):

it is very simple: in each integral we have to change variable, like below: \[\large {x^2} + 1 = t\] where t is the new variable

OpenStudy (anonymous):

Yeah Im going to use \[\frac{ x }{ (x^2+1)^2 } -\frac{x}{(x^2+1)^3}\] and split them into separate integrals and try doing u-substitution

OpenStudy (michele_laino):

yes! we have: \[\Large 2x\;dx = dt\]

OpenStudy (anonymous):

Ohh God, I solved it :D

OpenStudy (anonymous):

Thanks for all the help @Michele_Laino

OpenStudy (anonymous):

Now onto the next one ><

OpenStudy (michele_laino):

Thanks @OdinMW